# Dynamics of Pseudoentanglement

### v1

@@ IV.1 Motivation and setup
In this section, we study the problem of how pseudoentanglement spreads in space, focusing on the setting of a smaller pseudoentangled system placed in contact with a larger, trivial system. This problem is reminiscent of analogous settings in the dynamics of entanglement: e.g., the growth of entanglement after quantum quenches where thermalization may happen due to the spreading of entangled quasiparticles in space Calabrese and Cardy (2005, 2009), or the equilibration of quantum spin chains connected to thermal baths at their edges Prosen and Žnidarič (2009); Znidaric et al. (2011); Morningstar et al. (2022); Sels (2022).
@@
We consider a situation in which we are given a random subset-phase ensemble $\mathcal{E}_{0}=\{\ket{\psi_{S_{0},f_{0}}}:\ S_{0}\sim{\rm unif.},\ f_{0}\sim{\rm unif.}\}$ on a system $A$ of $N_{A}$ qubits. We take $|S_{0}|=K$ with $\omega({\rm poly}(N_{A}))<K<o(2^{N_{A}})$, so that the initial ensemble on $A$ is pseudoentangled. We then introduce a system $\bar{A}$ of $N-N_{A}$ ancillas initialized in the product state |$\ket{0}^{\otimes N-N_{A}}$, and form a larger system $A\bar{A}$ of $N$ qubits. This overall system is still in a subset-phase state, though not a uniformly-random one; the initial state indeed is of the form |$\ket{\psi_{S_{0},f_{0}}}\otimes\ket{0}^{\otimes N-N_{A}}$. we can straigthforwardly redefine $S_{0}$ as a subset of the computational basis of the extended system $\{0,1\}^{N}$ (rather than the original system $\{0,1\}^{N_{A}}$), by padding each bitstring with $N-N_{A}$ trailing zeros: $\mathbf{z}=(z_{1},\dots z_{N_{A}})\mapsto(z_{1},\dots z_{N_{A}},0,\dots 0)$. The phase function $f_{0}$ is never evaluated on any bitstrings outside the original ones, so it can be trivially extended to a uniformly random binary function on $\{0,1\}^{N}$. Thus after including the ancillas, we have a subset-phase ensemble with uniformly random phases, but highly-correlated subsets. These states are manifestly not pseudoentangled—one can efficiently distinguish them from Haar-random states by, e.g., measuring $\langle Z_{j}\rangle$ for any $j>N_{A}$.
@@
We then place the original system and ancillas into contact by allowing them to evolve together under automaton circuit dynamics, as discussed in Sec. III. The question we aim to address is if and how pseudoentanglement is achieved on the entire system of $N$ qubits over the course of the evolution. As the phases are already randomized, it is sufficient to focus on the subsets.
@@
A first apparent constraint is that the value of $K$ is conserved, even as we add the ancillas: even assuming that the subsets indeed equilibrate to the uniform distribution, pseudoentanglement requires $\omega({\rm poly}(N))<K<o(2^{N})$ (see Sec. II.1). Combining these constraints with the analogous ones from the initial state ($N\mapsto N_{A}$) yields the overall constraint [^6] Some examples of parameters satisfying these constraints are as follows: $N_{A}=N/2$, $K=\exp(N^{1/2})$; $N_{A}=N^{1/2}$, $K=\exp(N^{1/4})$; $N_{A}=(\log N)^{2}$, $K=N^{\log\log N}$. These examples are purely illustrative; in the following we will not assume anything about $N$, $N_{A}$, $K$ other than Eq. (20).
$\omega({\rm poly}(N))<K<o(2^{N_{A}}).$ (20)
@@
IV.2 All-to-all circuit model
Figure 4: All-to-all automaton circuit model for pseudoentanglement spreading. (a) Sketch of the circuit, showing a subset-phase initial state |$\ket{\psi_{S_{0},f_{0}}}$ on subsystem $A$ along with |$\ket{0}^{\otimes N-N_{A}}$ on the rest of the system, and a sequence of random globally-controlled NOT gates, $\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}$. Note that the controls come in two variations, open and solid circles, representing control on 0 and 1 respectively. (b) Sketch of the exact result Eq. (25) for the pseudothermalization transitions in this model. Consecutive pseudothermalization times $t_{\rm PTherm}^{(m)}$, $t_{\rm PTherm}^{(m+1)}$ are not sharply separated.
We start by analyzing a model with a high degree of mathematical tractability, that comes at the cost of all-to-all interactions and exponentially-long equilibration time scales. We emphasize that these latter aspects are not at all necessary to pseudothermalization in general, and are in fact absent from another model that we study next, in Sec. IV.3.
We consider a gate set made of globally-controlled NOT gates: $\mathcal{G}=\{\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}\}$, where $\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}$ flips qubit $i$ if and only if the remaining $N-1$ qubits are in computational basis state |$\ket{\mathbf{a}}$. There are $ND/2$ gates in $\mathcal{G}$: $N$ choices for the target $i$ and $22^{N-1}=D/$ 2 choices for the control string $\mathbf{a}$. The resulting circuit is illustrated in Fig. 4(a).
@@ IV.2 All-to-all circuit model
$(\mathbf{a}_{1:i-1},0,\mathbf{a}_{i+1:N})\xleftrightarrow{\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}}(\mathbf{a}_{1:i-1},1,\mathbf{a}_{i+1:N})$ (21)
@@
(we use the shorthand $\mathbf{a}_{i:j}=(a_{i},\dots a_{j})$). Clearly by a sequence of such neighbor transpositions one can transpose two arbitrary vertices [^7] Given a path from $\mathbf{x}$ to $\mathbf{y}$ on the hypercube, the sequence of neighbor transpositions along the path and back has the effect of transposing $\mathbf{x}$ and $\mathbf{y}$ while leaving all other sites fixed. $\mathbf{x}\leftrightarrow\mathbf{y}$; with these one can generate the whole permutation group $\mathcal{S}_{D}$. It follows that $\mathcal{G}$ induces an irreducible Markov chain on $\Sigma_{m}$ for all values of $m$. The Markov chains induced by $\mathcal{G}$ on the subset spaces $\{\Sigma_{m}\}$ are also clearly reversible ($\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}$ is its own inverse) and aperiodic (it is always possible to draw a value of $\mathbf{a},i$ such that $\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}$ acts trivially on all bitstrings in $S$, making the walk idle). By the discussion in Sec. III we have that the unique steady state of the Markov chain on $\Sigma_{K}$ is the uniform distribution $\pi_{K}$. Thus at late times this dynamics produces the random subset-phase state ensemble over the entire system, which in the relevant regime of $(K,N)$ is pseudoentangled.
@@
Having addressed the $t\to\infty$ limit of the dynamics, it remains to understand how pseudoentanglement is achieved as a function of time. To this end, we can use Proposition 1, which relates the $m$-copy distinguishability between the time-evolved state ensemble $\mathcal{E}_{t}$ and the Haar ensemble to the mixing of a Markov chain $\Phi_{m\leftarrow K}[p_{t}]$ on $\Sigma_{m}$.
@@
We note that the Markov chain on the permutation group $\mathcal{S}_{D}$ induced by this gate set $\mathcal{G}$ is precisely the interchange process (IP) on the hypercube $\mathbb{Z}_{2}^{N}$, reviewed in Sec. II.3: each gate picks an edge of the hypercube uniformly at random and transposes its two vertices. It follows that the Markov chains $\Phi_{m\leftarrow K}[p_{t}]$ over the spaces of subsets $\Sigma_{m}$ are exactly the exclusion processes $\textsf{EX}(m)$. These are well-studied problems for which we can rely on powerful known results.
@@
First of all, Aldous’ spectral gap conjecture Caputo et al. (2010), Eq. (10), states that the relaxation time of $\textsf{EX}(m)$ is equal, for all $m$, to the relaxation time of the simple random walk (RW). The simple random walk on the hypercube can be solved straightforwardly (see Appendix D.1), and one finds888 The exponential scaling in $N$ is due to the fact that each elementary transposition of hypercube vertices is exponentially unlikely to cause the random walker to hop: only $N$ gates out of $ND/2$ cause a nontrivial hop. The conventional result for a random walker who hops once per unit time is $t_{\rm rel}^{\textsf{RW}}=N/2$; rescaling this by the average number of gates between hops ($D/2$) yields Eq. (22) (see also more detailed derivation in Appendix D.1). $t_{\rm rel}^{\textsf{RW}}=ND/4$; therefore
@@
Secondly, the mixing time for $\textsf{EX}(m)$ was recently solved Hermon and Pymar (2020) and was found to scale as
@@
Due to Proposition 1, this mixing time is also the time taken to achieve $m$-copy pseudothermalization.
@@
This asymptotic separation between the scaling of relaxation and mixing times is a necessary condition for a cut-off phenomenon, reviewed in Sec. II.2. It was conjectured that in sufficiently “generic” models, such separation should also be a sufficient condition999 Note that the condition $t_{\rm rel}=o(t_{\rm mix})$ is not sufficient to get a cutoff phenomenon in total variation distance in general (on the other hand it is sufficient if the distance from equilibrium is measured with a $L^{p}$ norm, $p>1$ Chen and Saloff-Coste (2008)). It is possible to have a so-called pre-cutoff where multiple step-like transitions arise in the total variation distance. However, this usually requires some particular structure in the connectivity of the underlying graph (e.g. bottlenecks separating highly connected regions). . This cutoff effect would imply a sequence of dynamical transitions in our dynamics: for any moment $m\leq{\rm poly}(N)$, there is a time scale at which the ensemble of output states becomes indistinguishable from Haar-random at the $m$-copy level; this $m$-copy pseudothermalization time in this model is given by
$t_{\rm PTherm}^{(m)}\sim N2^{N}[\log(mN)\pm O(1)],$ (25)
@@
with the $\pm$ term denoting the temporal width of the dynamical crossover to equilibrium. This result is sketched in Fig. 4(b). We note that, while each pseudothermalization transition is sharp (in the sense of $t_{\rm rel}^{(m)}/t_{\rm mix}^{(m)}\to 0$ at large $N$), consecutive transitions $t_{\rm PTherm}^{(m)}$, $t_{\rm PTherm}^{(m+1)}$ are not sharply separated. This is due to the slow growth of $t_{\rm mix}^{(m)}$ with $m$, namely the fact that $[t_{\rm mix}^{(m+1)}-t_{\rm mix}^{(m)}]/t_{\rm rel}^{(m)}\sim\log\frac{m+1}{m}$ does not diverge (and in fact it goes to zero at large $m$).
@@
This exactly solvable model serves as a nice illustration of the idea of pseudothermalization, including the possible emergence of a cutoff phenomenon. At the same time, it has several drawbacks: the all-to-all interactions require complex couplings or compilation into deep local circuits; even worse, the resulting time scales for relaxation and mixing are exponentially long in system size. However, these limitations are artifacts of the model, and not intrinsic to the concept of pseudoentanglement or pseudothermalization. Below we introduce a local automaton gate set that does not suffer from these issues.
IV.3 Local circuit model
@@
Figure 5: Local automaton circuit model for the spreading of pseudoentanglement. (a) Circuit model, analogous to Fig. 4 but with local $\textsf{CC}X$ gates. (b) Schematic summary of bounds on the pseudothermalization time. Solid lines denote unconditional bounds while the dot-dashed line depends on Conjecture 1. The yellow area indicates allowed scalings. The lower bounds come from Eq. (30) (counting) and Eq. (31) (locality). The dashed-line upper bound comes from Eq. (29); the solid-line upper bound is Eq. (27). (c) Consecutive $m$-copy pseudothermalization transitions in this model may be overlapping (left) or sharply separated (right) in the thermodynamic limit: answering this question requires tighter bounds on the mixing time.
@@
In the interest of building a more physically realistic model of dynamics, we seek to restrict the gate set $\mathcal{G}$ to geometrically local gates on a $d$-dimensional array of qubits.
@@
The simplest automaton gate set comprises just local bit flips: $\mathcal{G}=\{X_{i}\}_{i=1}^{N}$. It is straightforward to see that the action of $\mathcal{G}$ on $\Sigma_{1}$ is irreducible (we can map any bitstring to any other via bitflips); however it is already not irreducible on $\Sigma_{2}$: given a set $S=\{\mathbf{z},\mathbf{z}^{\prime}\}$, the “relative coordinate” $\mathbf{z}\oplus\mathbf{z}^{\prime}$ is invariant under bit flips. Moving up in complexity, one might try two-qubit automaton gates, generated by NOTs and CNOTs; however, it is possible to show that these gates are also insufficient to achieve irreducibility on $\Sigma_{m}$ above $m=3$. Three-qubit automaton gates, on the other hand, can induce irreducible and aperiodic Markov chains on $\Sigma_{m}$ for all $m\leq D/4$, as we show in Appendix E.
@@
In particular, for the case of a one-dimensional qubit array, we consider the gate set $\mathcal{G}=\{u_{iab}=\textsf{C}_{i-1,a}\textsf{C}_{i+1,b}{X}_{i}\}$ where $a,b\in\{0,1\}$ label a control bitstring and $i\in[N]$ labels the target site (we take periodic boundary conditions, identifying $i=0$ and $i=N$), as shown in Fig. 5(a). The $u_{iab}$ gate flips qubit $i$ if and only if its left neighbor $i-1$ is in state $a$ and its right neighbor $i+1$ is in state $b$. There are $N$ choices for the target site $i$ and 4 choices for the control string $(a,b)$, giving a total of $4N$ gates in $\mathcal{G}$. Up to bit flips, each $u_{iab}$ is a Toffoli gate. It is easy to see that the gate set generates $\textsf{C}X$ gates, SWAP gates, and bit flip gates as well—it in fact generates arbitrary even permutations of the computational basiss, see Appendix E.
IV.3.1 First moment

### v3

@@ 4.1 Motivation and setup
In this section, we study the problem of how pseudoentanglement spreads in space, focusing on the setting of a smaller pseudoentangled system placed in contact with a larger, trivial system. This problem is reminiscent of analogous settings in the dynamics of entanglement: e.g., the growth of entanglement after quantum quenches where thermalization may happen due to the spreading of entangled quasiparticles in space calabrese_evolution_2005 ; calabrese_entanglement_2009 , or the equilibration of quantum spin chains connected to thermal baths at their edges Prosen_matrix_2009 ; znidaric_transport_2011 ; morningstar_avalanches_2022 ; sels_bath-induced_2022 .
@@
We consider a situation in which we are given a random subset-phase ensemble $\mathcal{E}_{0}=\{\ket{\psi_{S_{0},f_{0}}}:\ S_{0}\sim{\rm unif.},\ f_{0}\sim{\rm unif.}\}$ on a system $A$ of $N_{A}$ qubits. We take $|S_{0}|=K$ with $\omega({\rm poly}(N_{A}))<K<2^{o(N_{A})}$, so that the initial ensemble on $A$ is pseudoentangled. We then introduce a system $\bar{A}$ of $N-N_{A}$ ancillas initialized in the product state |$\ket{0}^{\otimes N-N_{A}}$, and form a larger system $A\bar{A}$ of $N$ qubits. This overall system is still in a subset-phase state, though not a uniformly-random one; the initial state indeed is of the form |$\ket{0}^{\otimes N-N_{A}}\otimes\ket{\psi_{S_{0},f_{0}}}$. We can straightforwardly redefine $S_{0}$ as a subset of the computational basis of the extended system $\{0,1\}^{N}$ (rather than the original system $\{0,1\}^{N_{A}}$), by padding each bitstring with $N-N_{A}$ leading zeros: $\mathbf{z}=(z_{1},\dots z_{N_{A}})\mapsto(0,\dots,0,z_{1},\dots z_{N_{A}})$. The phase function $f_{0}$ is never evaluated on any bitstrings outside the original ones, so it can be trivially extended to a uniformly random binary function on $\{0,1\}^{N}$. Thus after including the ancillas, we have a subset-phase ensemble with uniformly random phases, but highly-correlated subsets. These states are manifestly not pseudoentangled—one can efficiently distinguish them from Haar-random states by, e.g., measuring $\langle Z_{j}\rangle$ for any $j>N_{A}$.
@@
We then place the original system and ancillas into contact by allowing them to evolve together under automaton circuit dynamics, as discussed in Sec. 3. The question we aim to address is if and how pseudoentanglement is achieved on the entire system of $N$ qubits over the course of the evolution. As the phases are already randomized, it is sufficient to focus on the subsets.
@@
A first apparent constraint is that the value of $K$ is conserved, even as we add the ancillas: even assuming that the subsets indeed equilibrate to the uniform distribution, pseudoentanglement requires $\omega({\rm poly}(N))<K<2^{o(N)}$ (see Sec. 2.1). Combining these constraints with the analogous ones from the initial state ($N\mapsto N_{A}$) yields the overall constraint
$\omega({\rm poly}(N))<K<2^{o(N_{A})}.$ (21)
@@
4.2 All-to-all circuit model
Figure 4: Spreading of pseudoentanglement in an all-to-all automaton circuit model. (a) Sketch of the circuit, showing a subset-phase initial state |$\ket{\psi_{S_{0},f_{0}}}$ on subsystem $A$ along with |$\ket{0}^{\otimes N-N_{A}}$ on the rest of the system, and a sequence of random globally-controlled NOT gates, $\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}$. Note that the controls come in two variations, open and solid circles, representing control on 0 and 1 respectively. (b) Sketch of the exact result Eq. (26) for the distinguishability transitions in this model. Consecutive indistinguishability times $t_{\rm indist}^{(m)}$, $t_{\rm indist}^{(m+1)}$ are not sharply separated.
We start by analyzing a model with a high degree of mathematical tractability, that comes at the cost of all-to-all interactions and exponentially-long equilibration time scales. We emphasize that these latter aspects are not at all necessary to ensemble pseudothermalization in general, and are in fact absent from another model that we study next, in Sec. 4.3.
We consider a gate set made of globally-controlled NOT gates: $\mathcal{G}=\{\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}\}$, where $\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}$ flips qubit $i$ if and only if the remaining $N-1$ qubits are in computational basis state |$\ket{\mathbf{a}}$. There are $ND/2$ gates in $\mathcal{G}$: $N$ choices for the target $i$ and $22^{N-1}=D/$ 2 choices for the control string $\mathbf{a}$ ($D=2^{N}$ is the Hilbert space dimension). An example of the resulting circuit family is illustrated in Fig. 4(a).
@@ 4.2 All-to-all circuit model
$(\mathbf{a}_{1:i-1},0,\mathbf{a}_{i+1:N})\xleftrightarrow{\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}}(\mathbf{a}_{1:i-1},1,\mathbf{a}_{i+1:N}),$ (22)
@@
where we used the shorthand $\mathbf{a}_{i:j}=(a_{i},\dots a_{j})$. Clearly by a sequence of such neighbor transpositions one can transpose two arbitrary vertices [^12] Given a path from $\mathbf{x}$ to $\mathbf{y}$ on the hypercube, the sequence of neighbor transpositions along the path and back has the effect of transposing $\mathbf{x}$ and $\mathbf{y}$ while leaving all other sites fixed. $\mathbf{x}\leftrightarrow\mathbf{y}$; with these one can generate the whole permutation group $\mathcal{S}_{D}$. It follows that $\mathcal{G}$ induces an irreducible Markov chain on $\Sigma_{m}$ for all values of $m$. The Markov chains induced by $\mathcal{G}$ on the subset spaces $\{\Sigma_{m}\}$ are also clearly reversible ($\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}$ is its own inverse) and aperiodic (since $|S|=m\ll 2^{N}$, it is always possible to draw a control-target pair $(\mathbf{a},i)$ such that $\textsf{C}^{(N-1)}_{\mathbf{a}}{X}_{i}$ acts trivially on all bitstrings in $S$, making the walk idle). By the discussion in Sec. 3 we have that the unique steady state of the Markov chain on $\Sigma_{K}$ is the uniform distribution $\pi_{K}$. Thus at late times this dynamics produces the random subset-phase state ensemble over the entire system, which in the relevant regime of $(K,N)$ is pseudoentangled.
@@
Having addressed the $t\to\infty$ limit of the dynamics, it remains to understand how pseudoentanglement is achieved as a function of time. To this end, we can use Proposition 1, which relates the $m$-copy distinguishability between the quantum state ensembles, $\mathcal{E}_{p_{t}}$ and Haar, to the mixing of a Markov chain $\Phi_{m\leftarrow K}[p_{t}]$ on $\Sigma_{m}$, with $p_{t}$ the probability distribution over subsets at time $t$ during the dynamics.
@@
We note that the Markov chain on the permutation group $\mathcal{S}_{D}$ induced by this gate set $\mathcal{G}$ is precisely the interchange process (IP) on the hypercube $\mathbb{Z}_{2}^{N}$, reviewed in Sec. 2.3: each gate picks an edge of the hypercube uniformly at random and transposes its two vertices. It follows that the Markov chains $\Phi_{m\leftarrow K}[p_{t}]$ over the spaces of subsets $\Sigma_{m}$ are exactly the exclusion processes $\textsf{EX}(m)$. These are well-studied problems for which we can rely on powerful known results.
@@
First of all, Aldous’ spectral gap conjecture caputo_proof_2010 , Eq. (11), states that the relaxation time of $\textsf{EX}(m)$ is equal, for all $m$, to the relaxation time of the simple random walk (RW). The simple random walk on the hypercube can be solved straightforwardly (see Appendix D.1), and one finds131313 The exponential scaling in $N$ is due to the fact that each elementary transposition of hypercube vertices is exponentially unlikely to cause the random walker to hop: only $N$ gates out of $ND/2$ cause a nontrivial hop. The conventional result for a random walker who hops once per unit time is $t_{\rm rel}^{\textsf{RW}}=N/2$; rescaling this by the average number of gates between hops ($D/2$) yields Eq. (23) (see also more detailed derivation in Appendix D.1). $t_{\rm rel}^{\textsf{RW}}=ND/4$; therefore
@@
Secondly, the mixing time for $\textsf{EX}(m)$ was recently solved hermon_exclusion_2020 and was found to scale as
@@
Due to Proposition 1, this mixing time is also the time taken to achieve $m$-copy indistinguishability.
@@
This asymptotic separation between the scaling of relaxation and mixing times is a necessary condition for a cut-off phenomenon, reviewed in Sec. 2.2. It was conjectured that in sufficiently “generic” models, such separation should also be a sufficient condition141414 Note that the condition $t_{\rm rel}=o(t_{\rm mix})$ is not sufficient to get a cutoff phenomenon in total variation distance in general (on the other hand it is sufficient if the distance from equilibrium is measured with a $L^{p}$ norm, $p>1$ chen_cutoff_2008 ). It is possible to have a so-called pre-cutoff where multiple step-like transitions arise in the total variation distance. However, this usually requires some particular structure in the connectivity of the underlying graph. . This cutoff effect would imply a sequence of dynamical transitions in our dynamics: for any number of copies $m\leq O({\rm poly}(N))$, there is a time scale at which the ensemble of output states becomes indistinguishable from Haar-random at the $m$-copy level; this $m$-copy indistinguishability time in the present model is given by
$t_{\rm indist}^{(m)}\sim N2^{N}[\log(mN)\pm O(1)],$ (26)
@@
with the $\pm$ term denoting the temporal width of the dynamical crossover to equilibrium. This result is sketched in Fig. 4(b). We note that, while each distinguishability transition is sharp (in the sense of $t_{\rm rel}^{(m)}/t_{\rm mix}^{(m)}\to 0$ at large $N$), consecutive transitions $t_{\rm indist}^{(m)}$, $t_{\rm indist}^{(m+1)}$ are not sharply separated. This is due to the slow growth of $t_{\rm mix}^{(m)}$ with $m$, namely the fact that $[t_{\rm mix}^{(m+1)}-t_{\rm mix}^{(m)}]/t_{\rm rel}^{(m)}\sim\log\frac{m+1}{m}$ does not diverge (in fact it goes to zero at large $m$).
@@
This exactly solvable model serves as a nice illustration of the idea of ensemble pseudothermalization, including the possible emergence of a cutoff phenomenon. At the same time, it has several drawbacks: the all-to-all interactions require complex couplings or compilation into deep local circuits; even worse, the resulting time scales for relaxation and mixing are exponentially long in system size. However, these limitations are artifacts of the model, and not intrinsic to the concept of pseudoentanglement or ensemble pseudothermalization. Below we introduce a local automaton gate set that does not suffer from these issues.
4.3 Local circuit model
@@
Figure 5: Spreading of pseudoentanglement in a local automaton circuit model. (a) Circuit model, analogous to Fig. 4 but with geometrically local $\textsf{CC}X$ gates. (b) Schematic summary of bounds on the indistinguishability time. The $x$ axis qualitatively indicates asymptotic scalings of $m$ vs $N$ as $N\to\infty$. Solid lines denote unconditional lower bounds while the dot-dashed line in an upper bound that depends on Conjecture 1. Red regions are ruled out while the yellow region is allowed. The lower bounds come from Eq. (31) (counting) and Eq. (32) (locality). The upper bound comes from Eq. (30). (c) Consecutive $m$-copy indistinguishability transitions in this model may be overlapping (left) or sharply separated (right) in the thermodynamic limit: answering this question requires tighter bounds on the mixing time.
@@
To build a more physically realistic model of dynamics, we seek to restrict the gate set $\mathcal{G}$ to geometrically local gates on a $d$-dimensional array of qubits.
@@
The simplest local automaton gate set comprises just local bit flips: $\mathcal{G}=\{X_{i}\}_{i=1}^{N}$. It is straightforward to see that the Markov chain induced by $\mathcal{G}$ on $\Sigma_{1}$ is irreducible (we can map any bitstring to any other via bitflips); however it is already not irreducible on $\Sigma_{2}$: given a set $S=\{\mathbf{z},\mathbf{z}^{\prime}\}$, the “relative coordinate” $\mathbf{z}\oplus\mathbf{z}^{\prime}$ ($\oplus$ being sum modulo 2) is invariant under bit flips. Moving up in complexity, one might try two-qubit automaton gates, generated by NOTs and CNOTs; however, it is possible to show that these gates are also insufficient to achieve irreducibility on $\Sigma_{m}$ above $m=3$. Three-qubit automaton gates, on the other hand, can induce irreducible Markov chains on $\Sigma_{m}$ for all $m\leq D/4$, as we show in Appendix E.
@@
In particular, for the case of a one-dimensional qubit array, we consider the gate set $\mathcal{G}=\{u_{iab}=\textsf{C}_{i-1,a}\textsf{C}_{i+1,b}{X}_{i}\}$ where $a,b\in\{0,1\}$ label a control bitstring and $i\in[N]$ labels the target site (we take periodic boundary conditions, identifying $i=0$ and $i=N$), as shown in Fig. 5(a). The $u_{iab}$ gate flips qubit $i$ if and only if its left neighbor $i-1$ is in state $a$ and its right neighbor $i+1$ is in state $b$. There are $N$ choices for the target site $i$ and 4 choices for the control string $(a,b)$, giving a total of $4N$ gates in $\mathcal{G}$. Up to bit flips, each $u_{iab}$ is a Toffoli gate. It is easy to see that the gate set generates $\textsf{C}X$ gates, SWAP gates, and bit flip gates as well. In fact it generates arbitrary even permutations of the computational basis, see Appendix E.
4.3.1 First moment